OpenStudy (anonymous):

find three pairs of acute angles such that the sine of the first angle equals the cosine of the second angle could someone please explain how i can find these three acute angles?

OpenStudy (anonymous):

ok i can tell you the answer if you like, but there is a better idea look at the unit circle on the last page of the attached cheat sheet. the first coordinate is cosine, the second coordinate is sine. see if you can match them up for example in one you will see the first coordinate is \(\frac{\sqrt{3}}{2}\) and the second will be \(\frac{1}{2}\) and in another you will see the first coordinate is \(\frac{1}{2}\) and the second is \(\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

Should really memorize these, they come up all the time....

OpenStudy (anonymous):

@satellite73 im still confused.. would sqrt2/2,sqrt2/2 be one acute angle ? :S

OpenStudy (anonymous):

(sqrt2)/2, (sqrt2)/2

OpenStudy (anonymous):

Yes sin45 ¨= cos 45 = 1/sqrt2 = sqrt2/2

OpenStudy (anonymous):

its asking me to find 3 acute angles ... isnt that only one? :S

OpenStudy (anonymous):

Yes, you asked about it....